How do you solve this proportion two ways: x-1/3=x+1/5

Answer by Alan3354(69443)

How do you solve this proportion two ways: x-1/3=x+1/5
---------
If it's (x-1)/3 = (x+1)/5
-------
Cross multiply
5(x-1) = 3(x+1)
etc
=======================
(x-1)/3 = (x+1)/5
Multiply thru by the LCD 15
5(x-1) = 3(x+1)
Same result. Not sure if that's considered a 2nd way.